OpenStudy (anonymous):

Which description matches the transformations y = cos x undergoes to produce y = 2cos(-x) - 4?

OpenStudy (anonymous):

the -x part does nothing because cosine is even and so \[\cos(-x)=\cos(x)\]

OpenStudy (anonymous):

\[2\cos(x)\] has a amplitude 2, so goes form -2 to 2 instead of from -1 to 1 then -4 at the end moves the whole graph down 4 units, instead of going from -2 to 2 it goes from -6 to -2

OpenStudy (anonymous):

A. Horizontal compression by factor 2, vertical shift by 4 units up, then a reflection across the x-axis B. Reflection across the y-axis, vertical stretch by factor 2, then a vertical shift down 4 units C. Reflection across the y-axis, vertical shift of 2 units, horizontal shift right by 4 units D. Horizontal shift left 1 unit, then vertical shift down by 4 unit

OpenStudy (anonymous):

B but that is a very stupid answer

OpenStudy (anonymous):

since you cannot see the reflection across the y - axis!

OpenStudy (apoorvk):

cosx and cos(-x) are same. becaus edoesnt matter how you turn the angles on the plane, cos takes same values so graph: